You toss a fair coin 4 times. What is the probability that (round to 4 decimal places)
a) you get all Heads?
b) you get at least one Tail?

Answer by math_tutor2020(3835)


Part (a)

1/2 represents the probability of getting heads because there's one side we want out of 2 total.

The probability of getting four heads in a row is (1/2)^4 = 1/16 = 0.0625

Answer: 0.0625

==================================
Part (b)

When flipping four coins, you either could get all four heads or you could get at least one tail. Both events cannot happen at the same time. This means the two events are complementary.

Imagine throwing a dart at a map. If we define the events "hits land" and "hits water", then the two events are complementary. One or the other, but not both, must happen. Or going back to the coin flip, the events "heads" and "tails" are complementary since there only two sides to a coin.

Complementary probabilities always add to 1
A = event getting all four heads
B = event of getting at least one tail
P(A) + P(B) = 1
P(B) = 1 - P(A)
P(B) = 1 - 0.0625
P(B) = 0.9375

0.0625 was the probability of getting all four heads, so 0.9375 ​is the probability of anything else (getting one or more tails).

Side note: he decimal 0.9375 converts to the fraction 15/16

Answer: 0.9375